OpenStudy (anonymous):

convert polar equation into cartesian equation r=-4tan(theta)sec(theta) help?

OpenStudy (anonymous):

i've gotten as far as \[\frac{ -4\tan(\theta) }{ \cos(\theta) }=r\] \[rcos(\theta)=-4\tan(\theta)\] \[x=-4\tan(\theta)\]

OpenStudy (anonymous):

but now i'm stuck. can't figure how to get y=mx+b

OpenStudy (anonymous):

there are conversion formulas between polar and rectangular...

OpenStudy (anonymous):

which are? the one's i've got don't help me

OpenStudy (anonymous):

\(r^2=x^2+y^2\)

OpenStudy (anonymous):

that doesn't help me...

OpenStudy (anonymous):

\[y=rsin \theta\]

OpenStudy (anonymous):

i have both of those.

OpenStudy (anonymous):

\(x=r~cos~theta\)

OpenStudy (anonymous):

ok, yes, i have all of those. i can't use any of them.....................listing equations i already have isn't helping....

OpenStudy (anonymous):

of course you have to plug it all in the formula...

OpenStudy (anonymous):

what i mean to your given problem...

OpenStudy (anonymous):

i did. which left me with x=-4tan(theta)...like i said before. i'm stuck there. i need y=mx+b...

OpenStudy (anonymous):

can I say something?

OpenStudy (anonymous):

will go to that later, first we have to transform the equation in x and y only...

OpenStudy (anonymous):

yes @OOOPS ... :)

OpenStudy (anonymous):